OpenStudy (anonymous):

A concessionaire at a football game pays a vendor $44 per game for selling hot dogs at $3.13 each. (a) Write a linear function that describes the income I the vendor earns for the concessionaire during the game if the vendor sells h hot dogs. I(h) = (b) Find the income if the vendor sells 247 hot dogs. (Round your answer to the nearest cent.) $

OpenStudy (anonymous):

I guess by income they mean the profit. That starts -44 dollars and gets bigger by 3.13 for each hot dog. So the equation follows. l(h)=-44+3.13h l(247)=-44+3.13*247=729.11